### v2

# DiffPO: Diffusion-styled Preference Optimization for Efficient Inference-Time Alignment of Large Language Models

Abstract
Inference-time alignment provides an efficient alternative for aligning LLMs with humans. However, these approaches still face challenges, such as limited scalability due to policy-specific value functions and latency during the inference phase. In this paper, we propose a novel approach, Diffusion-styled Preference Optimization (DiffPO), which provides an efficient and policy-agnostic solution for aligning LLMs with humans. By directly performing alignment at sentence level, DiffPO avoids the time latency associated with token-level generation. Designed as a plug-and-play module, DiffPO can be seamlessly integrated with various base models to enhance their alignment. Extensive experiments on AlpacaEval 2, MT-bench, and HH-RLHF demonstrate that DiffPO achieves superior alignment performance across various settings, achieving a favorable trade-off between alignment quality and inference-time latency. Furthermore, DiffPO demonstrates model-agnostic scalability, significantly improving the performance of large models such as Llama-3-70B.
1 Introduction
The alignment of large language models (LLMs) with human preferences has recently emerged as a focal area of research [62, 53]. Prominent techniques such as Reinforcement Learning from Human Feedback (RLHF) [47] and Direct Preference Optimization (DPO) [50] have demonstrated substantial efficacy. However, these methods require the optimization of individual policies, posing challenges such as high consumption of training resources. Inference-time alignment [45, 27] provides an efficient alternative through direct adjustment of the model’s output distribution, thus avoiding the need for resource-intensive retraining. Despite its advantages, this approach still requires policy-specific value functions, limiting its scalability across different models. Additionally, the inference-time latency remains high, presenting further challenges to its practical deployment.
Figure 1: Comparison with Inference-Time Methods. Points closer to the top-right indicate a superior trade-off between performance and inference time.
Figure 2: Illustration of the DiffPO Framework. (a) The objective of LLM alignment is to adjust the output of LLMs to reflect human values and intentions. In this process, preferences are considered at the sentence level, focusing on aspects such as the style and format of the complete output. (b) We propose Diffusion-style Preference Optimization (DiffPO), which reconceptualizes the alignment process as a sentence-level denoising process, where the goal is to transform an unaligned sentence $\mathbf{y}^{(0)}$ into an aligned sentence $\mathbf{y}^{(T)}$ step by step. (c) Designed as a plug-and-play module, DiffPO can be directly integrated with the base model output and yield better alignment.
In this paper, we investigate an efficient and policy-agnostic preference optimization method. We begin by reconsidering the objective of aligning with humans [65, 53]. As illustrated in Fig. 2(a), the alignment process operates at the sentence level, focusing on adjusting key components of the generated content, such as style or format, to better reflect human intentions or values. Inspired by the global controllability of the diffusion process [38, 43], we propose Diffusion-styled Preference Optimization (DiffPO). DiffPO draws an analogy from the diffusion-based denoising process to model the iterative adjustment required for aligning human preferences, as shown in Fig. 2(b). By employing parallel decoding [51, 37], DiffPO directly predicts sentence-level transitions, thus avoiding the time latency associated with token-level generation. During the training phase, we optimize the DiffPO with an objective that maps generations with varying alignment levels to an aligned target, making it a policy-agnostic, plug-and-play module. The optimized DiffPO can then be seamlessly integrated with the output of the base model, enhancing its alignment level, as demonstrated in Fig. 2(c).
We evaluate the performance of DiffPO on several benchmark datasets, including AlpacaEval 2 [21], MT-bench [74], and HH-RLHF [5]. Empirical results demonstrate that DiffPO achieves superior alignment performance across various base models and settings. Compared to inference-time alignment techniques, DiffPO strikes an optimal trade-off between alignment performance and inference-time latency, as shown in Fig. 1. Additional experiments highlight the model-agnostic scalability of DiffPO across different base models. Specifically, DiffPO-9B significantly enhances the performance of models such as Llama-3-70B and GPT-4o, showcasing its capability to improve weak-to-strong supervision.
The advantages of DiffPO can be summarized as:
  • Model-agnostic. DiffPO is optimized to learn sentence-level refinement, independent of the specific base LLMs. This allows it to be applied across a variety of base LLMs. Furthermore, DiffPO does not require access to model parameters, which enhances its compatibility with API-based models and existing preference-aligned models.
  • Training and Inference Efficiency. As a post-inference alignment strategy, DiffPO adopts a one-for-all approach: it involves training one single DiffPO and applying it for all base models, thus significantly reducing the resource intensiveness associated with policy optimization. Moreover, by framing alignment as sentence-level prediction, DiffPO bypasses the time latency associated with token-level generation, thereby improving inference-time efficiency.
2 Method
2.1 Preliminaries: Large Language Models
Next-Token Prediction.
The text generation of autoregressive large language models (LLMs) with prompt $\mathbf{x}$ and response $\mathbf{y}$ can be modelled as a next-token prediction process. Given the input $\mathbf{x}$, The language model $\pi(\cdot|\mathbf{x})$ autoregressively maps from current tokens $(\mathbf{x},\mathbf{y}_{1:n-1})$ to a distribution over the next token $\mathbf{y}_{n}$. The maximum token, $N$, sets the length limit for LLM outputs, which conclude with an end-of-sentence (EoS) token $\mathbf{y}_{N}=$ EoS that ends the generation. The generated output $\mathbf{y}$ consists of predicted tokens $(\mathbf{y}_{1},\mathbf{y}_{2},...,\mathbf{y}_{N})$.
Alignment of LLMs.
During the alignment of LLMs, the objective is to optimize a language model $\pi_{\theta}$ that maximizes the user’s preference [15, 47, 50]:
$\max_{\pi_{\theta}}\mathbb{E}_{\begin{subarray}{c}x\sim D,\mathbf{y}\sim\pi_{\theta}(\mathbf{y}|\mathbf{x})\\ \mathbf{y}^{\prime}\sim\pi_{\mathrm{ref}}(\mathbf{y}|\mathbf{x})\end{subarray}}$ $[p(\mathbf{y}\succ\mathbf{y}^{\prime}|\mathbf{x})$
$-\beta D_{KL}(\pi_{\theta}\|\pi_{\mathrm{ref}})],$ (1)
where $p(\mathbf{y}\succ\mathbf{y}^{\prime}|\mathbf{x})$ represents the preference, i.e., the probability that $\mathbf{y}$ is preferred over $\mathbf{y}^{\prime}$ given the context $\mathbf{x}$, which can be generally represented by the reward function $r$. The parameter $\beta$ controls the deviation from the reference policy $\pi_{\mathrm{ref}}$, which generally corresponds to the SFT model.
Parallel Decoding of LLMs.
In comparison to next-token prediction, where token-level generation is performed sequentially to obtain a sentence, parallel decoding has demonstrated the capacity by enabling sentence-level generation and improving content quality [51, 37]. Concretely, supposing
$f(\mathbf{y}_{n},\mathbf{y}_{<n},\mathbf{x}):=\mathbf{y}_{n}-\arg\max_{\mathbf{y}}\pi(\mathbf{y}|\mathbf{y}_{<n},\mathbf{x}),$
parallel decoding re-frames the LLM inference process as solving a system of nonlinear equations w.r.t. all tokens in a sentence $\mathbf{y}_{n}$ for $n=1,\ldots,N$. It can be solved in a parallel and iterative way:
[…]
In this way, for one forward pass of the LLM at time $t$, we can obtain the next sentence $\mathbf{y}^{(t+1)}$ based on the previous one $\mathbf{y}^{(t)}$.
2.2 Diffusion-styled Preference Optimization
Motivation.
The goal of LLM alignment is to align the outputs of LLMs with human values or intentions [65]. In this process, preferences are defined at the sentence-level, focusing on the style or format of complete generated answers, as illustrated in Fig. 2(a). However, the generation of these responses occurs at the token level, following the next-token prediction pattern inherent in LLM modeling. This requires existing alignment techniques to optimize preferences (or rewards) at the token-level, which complicates the learning process [4, 75, 69]. This inconsistency prompts us to reconsider the formulation of the alignment process.
Reformulation.
Inspired by the potential benefits of the diffusion process in controllable text generation [24, 28, 67], we draw an analogy between the aligning LLMs and the diffusion process. Specifically, we propose Diffusion-styled Preference Optimization (DiffPO), which reconceptualizes alignment as a sentence-level denoising process. The denoising process $\pi$ gradually refines the initial unaligned output $\mathbf{y}^{(0)}$ by adjusting the format or style as a whole. This process ultimately produces the aligned output $\mathbf{y}^{(T)}$, as illustrated in Fig. 2(b). The sentence-level alignment process can be formulated as follows:
$\pi(\mathbf{y}^{(0:T)}):=p(\mathbf{y}^{(0)})\prod_{t=1}^{T}\pi(\mathbf{y}^{(t)}|\mathbf{y}^{(t-1)},\mathbf{x}),$ (3)
where $\mathbf{y}^{(0)}$ and $\mathbf{y}^{(T)}$ represent the initial unaligned and final aligned generations, respectively. The intermediate sequence $\mathbf{y}^{(1:T-1)}$ can be viewed as the unaligned generations progressively transitioning along the trajectory from $\mathbf{y}^{(0)}$ to $\mathbf{y}^{(T)}$.
Assuming the existence of a reward model $r(\mathbf{x},\mathbf{y})$, which captures how well the generated output $\mathbf{y}$ aligns with human preferences given the input $\mathbf{x}$, the goal is to optimize a DiffPO model $\pi_{\theta}$. This model learns to take a sentence as input and predict the next sentence with a higher reward, as illustrated in Fig. 2(c). The goal can be expressed as follows:
$\pi_{\theta}(\mathbf{y}^{(t)}|\mathbf{y}^{(t-1)},\mathbf{x})\propto p(\mathbf{y}^{(t-1)},\mathbf{x}){\text{exp}}(r(\mathbf{x},\mathbf{y}^{(t)})).$
By employing parallel decoding, the DiffPO model directly performs sentence-level predictions.
2.3 Consistency Optimization of DiffPO
Inspired by Consistency LLMs [35], we propose to consistently map any intermediate (unaligned) generation $\mathbf{y}^{(t)}$ to the aligned generation $\mathbf{y}^{(T)}$. We jointly optimize the DiffPO model $\pi_{\theta}$ with two losses: one aligns the intermediate generation with the aligned generation, and the other prevents the corruption of the autoregressive (AR) modeling in the base model, thereby maintaining the generation quality.
Consistency Loss.
For a prompt $\mathbf{x}$ with an unaligned generation $\mathbf{y}^{(t)}$, we directly guide the model to output $\mathbf{y}^{(T)}$ with $\mathbf{y}^{(t)}$ as the input by minimizing the following loss $L_{\mathrm{Con}}$=
$\mathbb{E}_{(\mathbf{x},\mathbf{y}^{(t)},\mathbf{y}^{(T)})\sim\mathcal{D}}\left[\sum_{i=1}^{N}\mathrm{KL}(\pi_{\theta^{-}}(\mathbf{y}_{<i}^{(T)},\mathbf{x})\|\pi_{\theta}(\mathbf{y}_{<i}^{(t)},\mathbf{x}))\right]$ (4)
where $\theta^{-}=\mathrm{stopgrad}(\theta)$ and $N$ denotes the length of generation. $\mathrm{KL}(\cdot\|\cdot)$ denotes the forward KL distance between two distributions.
AR Loss.
To prevent the corruption of the autoregressive (AR) modeling in the base model and maintain the generation quality, we incorporate the AR loss based on the generated sequence $\mathbf{y}^{(T)}$:
$L_{\mathrm{AR}}=\mathbb{E}_{(\mathbf{x},\mathbf{y}^{(T)})\sim\mathcal{D}}\left[-\sum_{i=1}^{N}\log\pi_{\theta}(\mathbf{y}^{(T)}_{i}|\mathbf{y}^{(T)}_{<i},\mathbf{x})\right].$ (5)
The total loss with weight $\omega$ is:
$L(\theta)=L_{\mathrm{AR}}+\omega L_{\mathrm{Con}}.$ (6)
2.4 The Objective of DiffPO within RLHF
In this section, we analyze the role of DiffPO in achieving the goal of RLHF. We start with the same RL objective as prior work, Eq. 2.1, under a general reward function $r^{*}$. Following prior work [49, 50], the optimal solution to the KL-constrained reward maximization objective in Eq. 2.1 takes the form: $r^{*}(\mathbf{x},\mathbf{y})=\beta\log\left(\frac{\pi^{*}(\mathbf{y}\mid\mathbf{x})}{\pi_{\mathrm{ref}}(\mathbf{y}\mid\mathbf{x})}\right)+\beta\log Z(\mathbf{x})$, where $Z(\mathbf{x})=\sum_{\mathbf{y}}\pi_{\mathrm{ref}}(\mathbf{y}|\mathbf{x})\exp\left(\frac{1}{\beta}r^{*}(\mathbf{x},\mathbf{y})\right)$ is the partition function. With Bradley-Terry model, we can represent the preference function as the difference of rewards for a preferred answer $\mathbf{y}_{w}$ and a dispreferred answer $\mathbf{y}_{l}$:
$p$ $(\mathbf{y}_{w}\succ\mathbf{y}_{l}|\mathbf{x})=\sigma(r^{*}(\mathbf{x},\mathbf{y}_{w})-r^{*}(\mathbf{x},\mathbf{y}_{l}))$
$=\sigma\left(\beta\log\frac{\pi^{*}(\mathbf{y}_{w}\mid\mathbf{x})}{\pi_{\mathrm{ref}}(\mathbf{y}_{w}\mid\mathbf{x})}-\beta\log\frac{\pi^{*}(\mathbf{y}_{l}\mid\mathbf{x})}{\pi_{\mathrm{ref}}(\mathbf{y}_{l}\mid\mathbf{x})}\right).$
Substitute by $\pi^{*}(\mathbf{y}\mid\mathbf{x})=\pi_{\mathrm{\textsc{{DiffPO}}}}(\mathbf{y}\mid\mathbf{y}^{\prime},x)$ $\pi_{\mathrm{ref}}(\mathbf{y}^{\prime}\mid\mathbf{x})$, we obtain $p(\mathbf{y}_{w}\succ\mathbf{y}_{l}|\mathbf{x})$ equals to
$\sigma\left(\beta\log\frac{\pi_{\mathrm{\textsc{{DiffPO}}}}(\mathbf{y}_{w}\mid\mathbf{y}_{l},\mathbf{x})}{\pi_{\mathrm{\textsc{{DiffPO}}}}(\mathbf{y}_{l}\mid\mathbf{y}_{l},\mathbf{x})}-\beta\log\frac{\pi_{\mathrm{ref}}(\mathbf{y}_{w}\mid\mathbf{x})}{\pi_{\mathrm{ref}}(\mathbf{y}_{l}\mid\mathbf{x})}\right).$ (7)
Note that the first term in Eq. 7 is optimized through the consistency loss in Eq. 4 by maximizing the probability of predicting $\mathbf{y}_{w}$. The second term depends only on $\mathbf{x}$, with $\pi_{\mathrm{ref}}$ remaining constant. Moreover, the deviation from the base policy can be easily controlled, since $\mathbf{y}_{w}$ is derived from $\mathbf{y}_{l}$.
In summary, the objective of DiffPO as defined in Eq. 6 aligns with the RLHF objective in Eq. 2.1. Furthermore, since $\pi_{\text{{{DiffPO}}}}$ is optimized independently from the base model $\pi_{\text{ref}}$, it can be deployed in a model-agnostic manner.
2.5 Practical Implementations
Generate Alignment Trajectories.
To implement DiffPO, we collect the alignment trajectory for each prompt, thereby forming an original training set $\mathcal{D}$. Specifically, for each prompt $\mathbf{x}$ from the UltraFeedback dataset [16], we generate $T$ responses using different base models. We then employ ArmoRM [61] reward model to score these responses. The response with the highest score is selected as $\mathbf{y}^{(T)}$. The remaining five responses are ranked based on their scores to form $\mathbf{y}^{(0:T-1)}$. $T$ is set to 6.
Training and Inference.
During the training phase, we initialize our aligning model $\pi_{\theta}$ using three backbones of varying sizes: Gemma-2-it-2B/9B, and Llama-3-8B-Instruct. The DiffPO model is optimized adhering to the optimization loss in Eq. 6 with parameters $N=256$ and $w=10^{3}$. Given the variable lengths of generations in $\mathcal{D}$, we standardize their lengths through padding or truncation. In the inference phase, the optimized model $\pi_{\theta}^{*}$ is employed to align responses from the vanilla generations produced by base models. Appendix 6.1 shows more implementation details.
Table 1: Comparison results with baseline methods. DiffPO achieves the superior alignment performance across all benchmarks, outperforming the training-based baselines (i.e., SimPO and DPO) in various settings. Notably, DiffPO requires only a single training session and is applicable to multiple base models. The best result is highlighted in bold, while the second-best result is highlighted with underline.
Method Llama-3-SFT (8B) Llama-3-Instruct (8B)
MT-bench AlpacaEval 2 HH-RLHF MT-bench AlpacaEval 2 HH-RLHF
GPT-4 LC (%) WR (%) Helpful Harmless GPT-4 LC (%) WR (%) Helpful Harmless
Base Model 6.21 22.09 20.81 0.59 0.91 6.78 36.83 42.12 0.67 0.93
w. DPO 6.59 29.84 36.77 0.68 0.89 6.90 47.20 53.56 0.74 0.92
w. SimPO 6.62 32.27 40.96 0.66 0.86 7.05 52.57 58.33 0.75 0.92
w. BPO 5.84 21.34 22.33 0.60 0.92 6.43 22.39 34.06 0.67 0.92
w. ARGS 6.14 9.06 13.97 0.49 0.86 6.84 31.83 34.74 0.64 0.89
w. BoN 6.79 35.14 32.26 0.62 0.92 6.89 45.10 49.94 0.67 0.92
w. Aligner 4.88 20.41 17.15 0.60 0.91 4.82 32.53 32.69 0.67 0.96
w. MetaAligner 4.46 19.81 18.23 0.52 0.89 4.50 20.75 19.08 0.52 0.91
w. DiffPO-8B 6.96 36.24 40.96 0.62 0.93 7.02 36.44 41.01 0.68 0.93
w. DiffPO-9B 7.45 49.72 54.23 0.71 0.98 7.40 55.84 61.88 0.72 0.98
Method Mistral-SFT (7B) Mistral-Instruct (7B)
MT-bench AlpacaEval 2 HH-RLHF MT-bench AlpacaEval 2 HH-RLHF
GPT-4 LC (%) WR (%) Helpful Harmless GPT-4 LC (%) WR (%) Helpful Harmless
Base Model 5.73 20.15 17.24 0.56 0.87 6.39 32.81 34.86 0.66 0.94
w. DPO 5.91 31.28 32.65 0.66 0.91 6.29 35.60 37.73 0.67 0.92
w. SimPO 6.17 31.16 33.72 0.63 0.86 6.36 35.78 40.21 0.67 0.93
w. BPO 5.55 18.23 17.23 0.64 0.92 5.99 19.61 27.49 0.66 0.93
w. ARGS 5.12 11.07 13.95 0.55 0.87 6.20 26.60 29.68 0.66 0.92
w. BoN 6.21 33.36 27.74 0.64 0.94 6.40 34.75 39.24 0.68 0.94
w. Aligner 4.27 18.27 15.53 0.60 0.95 4.42 28.88 30.30 0.66 0.93
w. MetaAligner 4.08 12.40 9.72 0.51 0.85 3.71 18.55 16.91 0.55 0.91
w. DiffPO-8B 6.87 34.42 40.08 0.62 0.88 7.04 35.92 40.70 0.68 0.92
w. DiffPO-9B 7.13 48.99 52.87 0.70 0.96 7.33 56.22 61.71 0.72 0.98
3 Experiment
3.1 Experiment Setup
Evaluation Benchmarks and Metrics.
We conduct our experiments using two widely recognized benchmarks for open-ended instruction-following: MT-Bench [74] and AlpacaEval 2 [21]. These benchmarks are designed to evaluate the conversational abilities of models across a diverse set of queries. AlpacaEval 2 includes 805 questions drawn from five distinct datasets, while MT-Bench covers eight categories and comprises a total of 80 questions. Additionally, we employ the HH-RLHF [5] datasets to assess how well the models’ generative capabilities align with human values, particularly emphasizing helpfulness and harmlessness. We adhere to each benchmark’s specific evaluation protocol to report scores. In AlpacaEval 2, we report both the raw win rate (WR) and the length-controlled win rate (LC), comparing performance against the GPT-4 model. In contrast, we present the average score for MT-Bench, also utilizing GPT-4 as the judge model. For HH-RLHF, we report scores that reflect the models’ helpfulness and harmlessness, as well as the overall score. These scores are measured using ArmoRM [61], a state-of-the-art reward model from RewardBench [36], designed to align with human preferences.
Table 2: Performance of DiffPO models. The results demonstrate that both DiffPO-2B and DiffPO-9B lead to performance improvements across different base models. DiffPO-9B enhances the performance of larger models, such as Qwen2.5-14B and 32B, as well as black-box GPT-4o, exhibiting a weak-to-strong improvement pattern. Furthermore, the results show that DiffPO can be effectively integrated with existing preference optimization methods, such as DPO and SimPO, further enhancing alignment performance.
Base Models MT-bench AlpacaEval 2 HH-RLHF
1-Turn 2-Turn Avg. LC (%) WR (%) Overall Helpful Harmless
Llama-3.2-1B-Instruct 5.32 5.13 5.25 15.57 19.09 0.0955 0.5978 0.9313
w. DiffPO-2B 6.97 5.97 6.47 39.42 44.20 0.1077 0.6948 0.9728
w. DiffPO-9B 7.56 6.95 7.30 50.70 56.08 0.1130 0.7059 0.9770
Llama-3.2-3B-Instruct 6.84 6.06 6.45 33.41 37.43 0.1037 0.6533 0.9183
w. DiffPO-2B 7.13 6.46 6.79 39.94 45.32 0.1069 0.6956 0.9682
w. DiffPO-9B 7.58 7.00 7.36 54.06 59.30 0.1132 0.7106 0.9875
Llama-3-8B-SFT+DPO 6.70 6.48 6.59 29.84 36.77 0.1044 0.6814 0.8900
w. DiffPO-9B 7.42 7.03 7.22 54.29 59.51 0.1134 0.7178 0.9765
Llama-3-8B-SFT+SimPO 6.63 6.61 6.62 32.27 40.96 0.1022 0.6640 0.8589
w. DiffPO-9B 7.59 7.08 7.42 55.66 60.67 0.1121 0.7156 0.9638
Llama-3-8B-it+DPO 6.75 7.05 6.90 47.20 53.56 0.1120 0.7387 0.9154
w. DiffPO-9B 7.79 6.98 7.39 58.56 63.60 0.1140 0.7211 0.9831
Llama-3-8B-it+SimPO 7.09 7.00 7.05 52.57 58.33 0.1143 0.7483 0.9182
w. DiffPO-9B 7.43 7.22 7.33 59.66 65.32 0.1142 0.7229 0.9756
Llama-3-70B-Instruct 7.41 7.59 7.5 46.14 51.12 0.1087 0.6928 0.9163
w. DiffPO-9B 8.23 7.28 7.75 58.18 62.34 0.1137 0.719 0.9757
Qwen2.5-3B-Instruct 6.73 5.59 6.16 35.52 40.42 0.1050 0.6802 0.9587
w. DiffPO-2B 7.06 6.26 6.66 42.63 47.83 0.1065 0.6973 0.9771
w. DiffPO-9B 7.58 7.24 7.41 55.71 61.43 0.1132 0.7106 0.9875
Qwen2.5-7B-Instruct 7.11 6.96 7.03 45.03 49.95 0.1095 0.6995 0.9442
w. DiffPO-2B 7.01 6.34 6.67 43.89 49.07 0.1074 0.7013 0.9659
w. DiffPO-9B 7.62 7.10 7.35 57.89 63.01 0.1117 0.7100 0.9445
Qwen2.5-14B-Instruct 7.24 6.71 6.98 51.60 57.10 0.1117 0.7100 0.9445
w. DiffPO-2B 7.08 6.33 6.71 43.70 48.76 0.1078 0.7017 0.9704
w. DiffPO-9B 7.62 7.35 7.48 55.13 60.65 0.1136 0.7185 0.9759
Qwen2.5-32B-Instruct 7.35 6.95 7.15 54.93 60.95 0.1132 0.7189 0.9594
w. DiffPO-9B 7.58 7.63 7.60 55.13 61.94 0.1143 0.7248 0.9797
GPT-4o (API) 7.40 7.47 7.43 53.64 62.01 0.1119 0.6974 0.9669
w. DiffPO-9B 7.66 7.37 7.51 58.91 64.30 0.1129 0.7167 0.9893
Figure 3: Comparison of Inference-Time Efficiency. We compare DiffPO with existing inference-time alignment techniques, evaluating both alignment performance and execution time. Points located closer to the top-right corner indicate a better trade-off. When considering both aspects, DiffPO demonstrates a surpassing performance-efficiency trade-off on all three datasets.
Baselines.
We compare DiffPO with two primary categories of offline preference optimization methods. In the category of training-based methods: Direct Preference Optimization (DPO) [50] reparameterizes reward functions to simplify and stabilize the preference learning process. SimPO [44] utilizes the average log probability of a sequence as an implicit reward, aligning more closely with model generation. For training-free methods: Black-Box Prompt Optimization (BPO) [14] adapts user prompts to better align with LLMs’ input comprehension, achieving user intents optimally without altering LLM parameters. ARGS [34] integrates alignment into the decoding process through reward-guided search, eliminating the need for costly RL training. Best-of-N sampling (BoN) [46] samples N times and selects the highest-scoring sample based on the reward model, with N set to 4 in our experiments using ArmoRM [61] as the reward model. Furthermore, Aligner [31] and MetaAligner [63] employ an additional model to learn corrective residuals between preferred and dispreferred responses to refine model generation.
Table 3: Performance Under Hybrid Decoding. We segment the vanilla generation into blocks of varying sizes and sequentially apply DiffPO-8B to each block. This approach allows DiffPO decoding to be parallel within blocks and auto-regressive between blocks. Hybrid decoding significantly reduces the decoding time, indicating a feasible trade-off for performance.
Block Size 16 32 64 128 256
MT (GPT-4) 6.86 6.96 6.84 6.81 6.77
Time (s) 1080 1012 1390 1520 1937
AE2 (LC) 33.52 33.54 33.46 32.98 36.24
Time (s) 3712 3471 4614 5510 7520
HH (Avg.) 0.7742 0.7749 0.7743 0.7741 0.7761
Time (s) 1564 1551 1620 1816 2684
Base Models and Inference Settings.
We perform preference optimization primarily on two model families: Llama-3-8B [3] and Mistral-7B [32], under two configurations: SFT and Instruct. In the SFT configuration, we utilize open-source models from SimPO [44] that follow Zephyr [60] to train the base models (i.e., meta-llama/Meta-Llama-3-8B) on the UltraChat-200k [20] dataset to derive an SFT model. For the Instruct configuration, we employ an off-the-shelf instruction-tuned models (i.e., meta-llama/Meta-Llama-3-8B-Instruct). To further validate scalability, we conduct additional experiments using the Llama-3.2 series, Qwen-2.5 series [59], and GPT-4o [1] as the base models.
During the inference phase of DiffPO, we initially generate responses using the base models. For each benchmark. In AlpacaEval 2 and HH-RLHF, we employ a sampling decoding strategy with a temperature setting of 0.7. For MT-Bench, we adhere to the official decoding configuration, which specifies varying temperatures for different categories. In our primary experiments, we set the maximum token generation length to 256. Results for experiments conducted at various lengths are provided in Tab. 4. Subsequently, the responses generated by the base models are aligned using the trained DiffPO. For the main results, parallel decoding is executed with a block size of 256.
3.2 Experiment Results
DiffPO significantly outperforms existing preference optimization methods.
As shown in Table 1, while all preference optimization algorithms improve performance over the base model, DiffPO achieves the best overall performance across all benchmarks and settings. These consistent and significant improvements underscore the robustness and effectiveness of DiffPO. Notably, DiffPO outperforms the training-based baselines (i.e., SimPO and DPO) across various settings, despite requiring only a single training session of DiffPO model and being capable of enhancing the performance of multiple base models.
Table 4: Scaling towards Longer Generation Lengths. We evaluate the performance of DiffPO  under various maximum length settings. When the same optimized DiffPO-8B and 9B is applied to these outputs, consistently enhanced performance demonstrates DiffPO’s robust scaling capabilities.
Generation Length 256 512 1,024 2,048
Llama-3-SFT 6.21 6.61 6.76 6.71
w. DiffPO-8B ($\Delta$) +0.75 +0.81 +0.93 +1.05
w. DiffPO-9B ($\Delta$) +1.24 +1.64 +1.48 +0.50
Llama-3-Instruct 6.78 7.87 7.99 8.00
w. DiffPO-8B ($\Delta$) +0.24 -0.12 +0.01 +0.02
w. DiffPO-9B ($\Delta$) +0.62 +0.68 +0.34 +0.62
Mistral-SFT 5.73 6.42 6.50 6.36
w. DiffPO-8B ($\Delta$) +1.14 +1.16 +1.25 +1.51
w. DiffPO-9B ($\Delta$) +1.40 +1.47 +1.71 +1.86
Mistral-Instruct 6.39 7.47 7.68 7.64
w. DiffPO-8B ($\Delta$) +0.65 +0.06 +0.13 +0.17
w. DiffPO-9B ($\Delta$) +0.94 +0.59 +0.58 +0.78
Table 5: Loss and Hyperparameter Ablation. The results indicate that applying DiffPO-9B to the base model (i.e., gemma-2-9B) yields outperforming performance than vanilla decoding of any single models, demonstrating the effectiveness of DiffPO strategy.
Loss MT AE2 HH
GPT-4 LC (%) Avg.
Vanilla Decoding
Gemma-2-9B-it 7.37 54.15 0.8265
$L_{\mathrm{AR}}$ 7.42 58.19 0.8279
Gemma-2-9B-it w. DiffPO-9B
$L_{\mathrm{AR}}$ only 7.41 58.01 0.8301
$10\times L_{\mathrm{Con}}+L_{\mathrm{AR}}$ 7.56 59.27 0.8389
$100\times L_{\mathrm{Con}}+L_{\mathrm{AR}}$ 7.55 59.43 0.8435
$1,000\times L_{\mathrm{Con}}+L_{\mathrm{AR}}$ 7.60 60.56 0.8438
DiffPO consistently improves the performance of base models of various sizes.
We report the performance of DiffPO-2B and DiffPO-9B on base models of various sizes, with the results presented in Table 2. The results demonstrate that both DiffPO-2B and DiffPO-9B lead to performance improvements across different base models. However, the performance gain of DiffPO-2B is limited, showing notable improvements primarily for smaller models. In contrast, DiffPO-9B enhances the performance of larger models, such as Qwen2.5-14B and 32B, as well as black-box GPT-4, exhibiting a weak-to-strong improvement pattern. Furthermore, the results show that DiffPO can be effectively integrated with existing preference optimization methods, such as DPO and SimPO, further enhancing alignment performance. These results underscore the scalability of DiffPO.
DiffPO achieves a surpassing performance-efficiency trade-off.
We compare DiffPO with existing inference-time alignment techniques, evaluating both alignment performance and execution time. The results are illustrated in Fig. 3, with the execution time measured on a single NVIDIA A100 80GB GPU. Points located closer to the top-right corner indicate a more favorable Pareto frontier. BoN and MetaAligner achieves commendable alignment performance and inference time respectively. However, when considering both aspects, DiffPO demonstrates a surpassing performance-efficiency trade-off on all three datasets. The experiments are conducted on Llama-3-SFT.
3.3 Analysis
Performance Under Hybrid Decoding.
We investigate the hybrid decoding strategy of DiffPO, with results provided in Tab 3. We segment the vanilla generation, which has a maximum length of 256, into blocks of varying sizes and sequentially apply DiffPO-8B to each block. This approach allows DiffPO decoding to be parallel within blocks and auto-regressive between blocks. It can be observed that hybrid decoding significantly reduces the decoding time, with optimal efficiency achieved at a block size of 32. On the other hand, performance is enhanced when the block size is set to 256, which corresponds to purely parallel decoding, indicating a feasible trade-off. The experiments are conducted on Llama-3-SFT.
Scaling towards Longer Generation Lengths.
We validate the scalability of the DiffPO model in response to increasing generation lengths, with results presented in Tab 4. Using base models, we generate outputs on MT-Bench under various maximum length settings and observe a positive correlation between increased text length and higher scores. Subsequently, the same optimized DiffPO-8B and 9B is applied to these outputs using the hybrid decoding strategy described in the previous section. This approach consistently yields enhanced alignment performance, demonstrating DiffPO’s robust scaling capabilities towards longer generation lengths.
Loss and Hyperparameter Ablation.
We evaluate the effectiveness of the training loss of DiffPO in Section 2.3 and the inference strategy in Section 2.5. The results are presented in Table 5. We report on two decoding strategies: vanilla decoding of a single model and DiffPO decoding, which applies the optimized DiffPO-9B on the output of the base model. The findings indicate that applying DiffPO to the base model achieves performance superior to that of single models alone, thus demonstrating the effectiveness of the DiffPO strategy. Furthermore, we report the results of an ablation study on the hyperparameter $w$ in Eq. 6. When using DiffPO decoding, employing $L_{\mathrm{Con}}$ with larger values of $w$ lead to a more pronounced improvement in performance.
4 Conclusion
This paper introduces a novel inference-time alignment framework for large language models, DiffPO. DiffPO achieves alignment at the sentence level to better model human preferences, drawing inspiration from the denoising process. DiffPO outperforms both strong training-based and inference-time alignment techniques in terms of alignment performance and inference speed. Experiments scaling DiffPO from 2B to 9B parameters, expanding the base model from 1B to 70B, and increasing the context length from 256 to 2,048 demonstrate that DiffPO is a robust and scalable framework for LLM alignment.
Limitations
We acknowledge the presence of certain limitations. While DiffPO has demonstrated a superior trade-off between performance and inference-time cost, it still introduces additional inference latency due to the need for an extra model for alignment. Moreover, we observe that the performance of DiffPO scales with its size, which presents challenges for cost-effectiveness during deployment. Additionally, despite the empirical success and intuitive motivation behind DiffPO, a more rigorous theoretical analysis is required to fully understand its effectiveness. Future work could explore how to combine the diffusion process (i.e., the denoising process) with the alignment task more effectively. This paper draws insights from the analogy between the denoising process and alignment. We hope our findings will facilitate future exploration of existing successful techniques in the natural language processing domain.
Potential Risks
As an inference-time alignment technique, DiffPO aims to develop AI assistants that align with positive human intentions and social values. However, there is a potential risk that DiffPO could be misused to align with harmful or negative values. We strongly oppose any such misuse, as it could hinder human progress, and advocate for the responsible and ethical use of DiffPO.
Supplementary Material
The supplementary material is structured as follows:
  • Related works in Section 5.
  • Additional experiment setup and results in Section 6.1.
  • Additional analysis and ablation studies in Section 7.
5 Related Works
5.1 Align LLM with Human Preference.
Large Language Models (LLMs) shows impressive performance in many tasks [54, 71, 55, 72, 9]. A prominent approach to learning from human preferences is RLHF [47, 58, 15, 5, 13]. In this framework, a reward model is first trained, followed by the training of a bandit policy using Proximal Policy Optimization (PPO) [52]. Recent advancements such as direct preference optimization (DPO) [50, 44, 22, 73] optimize the bandit policy directly from human preferences, bypassing the need for a reward model. These approaches are simpler to implement and require fewer computational resources. Inference-time approaches, on the other hand, achieve alignment by customizing the output of large language models (LLMs) during the decoding phase, without the need for parameter optimization. This results in enhanced flexibility and efficiency [34, 45]. One representative method treats the text-generation process as a search problem, guided by external rewards [30, 27, 10]. Another category of methods focuses on learning to refine the generated text [39, 31, 63].
Token and Sentence-level.
Existing training-based or inference-time alignment approaches typically rely on token-level rewards, while human preferences are generally provided and defined at the sentence level [40, 2, 69]. To address this discrepancy, some works [41, 64, 69] leverage token-wise or step-wise information to improve alignment performance. In contrast, this paper proposes modeling alignment as a sentence-level denoising process. We introduce a model-agnostic, inference-time alignment method, and our empirical results demonstrate its superiority in both performance and efficiency.
5.2 Parallel Decoding and Diffusion Process.
Parallel Decoding of LLMs
Parallel decoding has been increasingly utilized and developed in recent research to accelerate the inference processes of large language models (LLMs). One line of research, including works by Leviathan et al. [37], Chen et al. [11], focuses on speculative decoding. These techniques enhance LLM decoding speed by employing a smaller draft model to predict the outputs, which are then verified in parallel by a larger target model. Another research trajectory explores parallel decoding strategies that do not rely on a draft model. Methods such as conditioning on “look-ahead” tokens or employing Jacobi iterations have been investigated by Santilli et al. [51], Fu et al. [23]. These approaches allow the target model to produce several tokens simultaneously, aiming for rapid convergence to a fixed point on a Jacobi trajectory. CLLMs [57] develop a novel approach, refining the target LLM to consistently predict the fixed point from any given state.
Table 6: Comparison results of DiffPO models. The experiments are conducted on base models of Qwen-2.5-7B and 14B. It shows that DiffPO consistently achieves superior performance across various base models.
Base Models MT-bench AlpacaEval 2 HH-RLHF
1-Turn 2-Turn Avg. LC (%) WR (%) Overall Helpful Harmless
Qwen2.5-7B-Instruct 7.11 6.96 7.03 45.03 49.95 0.1095 0.6995 0.9442
w. DPO 7.37 6.96 7.17 50.55 55.00 0.1109 0.7061 0.9460
w. SimPO 7.41 6.98 7.20 48.76 52.75 0.1100 0.7047 0.9387
w. BPO 6.90 6.16 6.53 29.44 39.85 0.1086 0.6765 0.9418
w. BoN 7.50 7.10 7.30 50.42 55.43 0.1159 0.7066 0.9440
w. Aligner 6.24 3.76 5.00 42.15 45.82 0.1088 0.6993 0.9438
w. MetaAligner 6.41 5.13 5.77 36.58 38.45 0.0995 0.6966 0.9422
w. DiffPO-2B 7.01 6.34 6.67 43.89 49.07 0.1074 0.7013 0.9659
w. DiffPO-9B 7.62 7.10 7.35 57.89 63.01 0.1117 0.7100 0.9445
Qwen2.5-14B-Instruct 7.24 6.71 6.98 51.60 57.10 0.1117 0.7100 0.9445
w. BPO 7.21 6.82 7.02 37.02 47.51 0.1005 0.6853 0.9323
w. BoN 7.49 6.98 7.24 54.92 59.02 0.1182 0.7163 0.9485
w. Aligner 6.14 4.11 5.13 45.14 47.24 0.1114 0.7099 0.9438
w. MetaAligner 6.24 5.73 5.99 41.25 43.23 0.1092 0.7074 0.9375
w. DiffPO-2B 7.08 6.33 6.71 43.70 48.76 0.1078 0.7017 0.9704
w. DiffPO-9B 7.62 7.35 7.48 55.13 60.65 0.1136 0.7185 0.9759
Text Diffusion Models
Diffusion models have demonstrated significant diversity and controllability in visual generation [29, 56, 19, 12, 8, 7, 33, 6, 17, 18, 48, 68]. Recently, these models have been extended to text generation, as evidenced by the works of [38, 24, 42]. In essence, diffusion models execute a multi-step denoising process that progressively transforms random noise into a coherent data sample. In the context of text, diffusion models can be considered an evolution of traditional iterative Non-Autoregressive models, as described by Gong et al. (2022). These models have demonstrated the ability to match or surpass Autoregressive (AR) models in terms of text perplexity [28, 26], diversity [25, 70], and various sequence-to-sequence tasks [67, 66].
Connection with DiffPO
In this paper, we are motivated by the goal of aligning Large Language Models (LLMs) with human values or intentions, as outlined in [65]. We define preferences at the sentence-level, focusing on the style or format of complete answers generated by the LLMs. If we consider each iteration of parallel decoding as a transition between states, this bears a formal resemblance to discrete diffusion models. In DiffPO, we leverage parallel decoding to implement sentence-level denoising, thereby enhancing the modeling of the alignment process.
The development of DiffPO is also inspired by Consistency Models [57] and CLLMs [35]. Consistency models address the limitation of the slow iterative sampling process by mapping any point along the probability flow ODE of the diffusion process back to the original point in a single step. CLLMs propose accelerating LLM inference by mapping the intermediate process of LLM parallel decoding to the final process. Similar to these works, we optimize DiffPO with consistency loss, thus enabling model-agnostic alignment.
6 Experiment
6.1 Experimental Setups
Training Details.
As for the training set, we collect 6 generations from 6 base models (i.e., Llama-3-8B-Instruct, Llama-3-8B-SFT, Mistral-7B-SFT, Mistral-7B-Instruct, Gemma-2-2B-Instruct, Gemma-2-9B-Instruct). We then employ ArmoRM [61] to score these responses. The response with the highest score is selected as $\mathbf{y}^{(T)}$. The remaining five responses are ranked according to their scores to serve as $\mathbf{y}^{(0:T-1)}$. In the training process, at each iteration, we randomly sample $\mathbf{y}^{t}$ from $\mathbf{y}^{(0:T-1)}$ for optimization. We train DiffPO models using the following hyperparameters: a learning rate of 1e-9, a batch size of 1 and gradient accumulation steps of 4, a max sequence length of 1024, and a cosine learning rate schedule with 3% warmup steps for 1 epoch. All the models are trained with an Adam optimizer. All the training experiments in this paper were conducted on 8×A100 GPUs.
Evaluation Details.
For the MT-bench, we use GPT-4 as the judge model, following the default settings. The scores are based on a single-answer rating scale from 1 to 10. For AlpacaEval, we use GPT-4 Turbo as the judge model, which performs pairwise comparison of responses generated by GPT-4, each with the same maximum length. For HH-RLHF, we use ArmoRM for single-answer rating and report the overall score, along with the “helpful” and “harmless” scores, which are provided in dimensions 9 and 10, respectively.
Baseline Details.
Implementation details for different baselines are as follows:
Figure 4: Illustration of the Speedup of DiffPO.
6.2 Experimental Results
DiffPO significantly outperforms existing preference optimization methods.
We provided additional comparison with baselines, with results presented in Table 6. The experiments are conducted on base models of Qwen-2.5-7B and 14B. While all preference optimization algorithms improve performance over the base model, DiffPO achieves the best overall performance across all benchmarks and settings. These consistent and significant improvements underscore the robustness and effectiveness of DiffPO. Notably, DiffPO outperforms the training-based baselines (i.e., SimPO and DPO) across various settings, despite requiring only a single training session of DiffPO model and being capable of enhancing the performance of multiple base models.
7 Analysis
Table 7: Loss and Hyperparameter Ablation. We report the results of vanilla decoding from the base model and the optimized DiffPO model. The results indicate that applying DiffPO to the base model yields outperforming performance than single models, demonstrating the effectiveness of DiffPO strategy.
Loss MT AE2 HH
GPT-4 LC (%) Avg.
Vanilla Decoding
Llama-3-Instruct 6.78 36.83 0.7985
$L_{\mathrm{AR}}$ 6.90 36.35 0.7971
Llama-3-it w. DiffPO-8B
$L_{\mathrm{AR}}$ only 6.75 35.84 0.7968
$10\times L_{\mathrm{Con}}+L_{\mathrm{AR}}$ 6.85 35.96 0.7997
$100\times L_{\mathrm{Con}}+L_{\mathrm{AR}}$ 6.86 35.92 0.7998
$1,000\times L_{\mathrm{Con}}+L_{\mathrm{AR}}$ 7.02 36.44 0.7998
7.1 Loss and Hyperparameter Ablation.
We supplement the ablation results for DiffPO-8B, presented in Table 7. We report on two decoding strategies: vanilla decoding of the single base model and the optimized DiffPO-8B, and DiffPO decoding, which applies the optimized DiffPO-8B on the output of the base model. The findings indicate that applying DiffPO to the base model achieves performance superior to that of single models alone, thus demonstrating the effectiveness of the DiffPO strategy. Furthermore, we report the results of an ablation study on the hyperparameter $w$ in Eq. 6. When using DiffPO decoding, employing $L_{\mathrm{Con}}$ with larger values of $w$ lead to a more pronounced improvement in performance.
7.2 Illustration of the Speed-up of DiffPO
As shown in Figure 4, AR decoding (e.g., Aligner [31]) typically generates only one aligned token per iteration. In contrast, DiffPO enables the skipping of satisfied tokens, thereby avoiding the time latency associated with token-level generation. As a result, DiffPOcan predict the modified subsequence in 3 iterations, achieving the same result as 11 iterations of AR decoding.